A householder uses panels of solar cells to generate electricity for his home. The solar cells are titled to receive the maximum energy input from the sun.

use the data in the table to describe how the average energy input to the solar cells depends on the angle of tilt.


1
Expert's answer
2021-01-31T07:56:37-0500

We can notice that the maximum power output occurs when the panels are "\theta=90°" to the sunlight, which means that


"P(\theta)=kP_S\text{ sin}\theta,"

where

"k" - a coefficient,

"P_S" - Power from the Sun.